# Nonlinear dynamics in Horndeski gravity: a renormalized approach to effective gravitational coupling

### v1

@@ 1 Introduction
The late-time accelerated expansion of the Universe remains a central puzzle in modern cosmology. While the standard $\Lambda$CDM model associates this phenomenon to a cosmological constant, many cosmologists have focused on the possibility that gravity itself could be modified on large scales. Among the broad classes of modified theories [1, 2], Horndeski’s theory [3, 4, 5] and its generalizations stand out as the most general scalar-tensor theory yielding second-order equations of motion and thus free of Ostrogradski-type instabilities, which arises when higher-order time derivatives introduce extra, potentially ghost-like degrees of freedom. Therefore these theories bring higher order derivatives in a consistent way. Indeed, the Galileon theory has the remarkable property that certain higher-derivative operators remain exactly non-renormalized by quantum corrections [6, 7]. Essentially, the shift symmetry forces loop corrections to vanish or be re-expressible in terms of lower-derivative terms. Unfortunately, once the Galileon field is coupled to gravity, this exact symmetry is broken and so one loses the strict non-renormalization. However, there are special subsets of these theories that preserve a remnant of the Galileon symmetry and the quantum corrections to the higher-derivative terms may remain suppressed [8, 9], preserving some of the non-renormalization arguments at least to some order or around special backgrounds.
@@
Early investigations of how Vainshtein screening impacts structure formation were carried out, for instance, in the DGP scenario [15], where a non-linear perturbative approach was developed suited to scales where screening becomes important. This paper generalizes the methods of [15] to a class of Horndeski-like models endowed with Vainshtein screening. Specifically, we formulate a renormalized perturbation theory framework designed to capture both the large-scale, linearized modifications to gravity and the small-scale screening behavior in a self-consistent way.
@@ 2 Viable Horndeski’s theories
where $\phi$ is the scalar field and $X=-(\partial_{\mu}\phi)^{2}/2$ is the kinetic term. The functions $K$ and $G$ are arbitrary.
@@ 4 Averaged effective gravitational constant
$\Phi(t,\mathbf{p})$ $=\int{\textnormal{d}}^{3}\mathbf{p}^{\prime}\left\langle\frac{\mathcal{D}\Phi(t,\boldsymbol{p})}{\mathcal{D}\delta_{m}(t,\boldsymbol{p}^{\prime})}\right\rangle\delta_{m}(t,\boldsymbol{p}^{\prime})+\frac{1}{2!}\int{\textnormal{d}}^{3}\mathbf{p}_{1}{\textnormal{d}}^{3}\mathbf{p}_{2}\left\langle\frac{\mathcal{D}^{2}\Phi(t,\boldsymbol{p})}{\mathcal{D}\delta_{m}(t,\boldsymbol{p}_{1})\mathcal{D}\delta_{m}(t,\boldsymbol{p}_{2})}\right\rangle\delta_{m}(t,\boldsymbol{p}_{1})\delta_{m}(t,\boldsymbol{p}_{2})+\ldots$ (39)
where $\mathcal{D}$ stands for a functional derivative. But because of translation invariance, we should have
@@ 7 Master equation
Going back to our problem, using the redefinition (76) and the nonlinear Poissson equation (3), we obtain
@@ 8 Renormalized power spectrum and bispectrum
The equations (86, 87) can be solved perturbatively to correct the assumed power spectrum and bispectrum. The algorithm is the following
@@
    Step 1: Start with an initial power spectrum $P(k)$ and bispectrum $B\left(k_{1},k_{2},k_{3}\right)$.
@@
    Step 3: Insert this back into the fluid equations (86,87,88) to obtain corrected $P(k)$ and $B\left(k_{1},k_{2},k_{3}\right)$.
@@
and $D_{1}=D_{0}+D_{2}$.
Given a renormalized effective gravitational constant, obtained at previous steps, we can in principle evolve easily eqs.(92,99,100) and consequently obtain the corrections to the power spectrum and the bispectrum which will be used in eq.(82) to correct the gravitational constant. These additional steps will be performed in a future work while we will address the construction of the renormalized gravitational constant in the next sections.
@@ 9 The model
To proceed with the analysis, we must adopt a specific model. As noted earlier, the coupling function $F(\phi)$ is not screened by the Vainshtein mechanism. To avoid fine-tuning, we assume $F(\phi)=1$. Rather than selecting a particular Lagrangian by specifying the remaining functions $K(\phi,X)$ and $G(\phi,X)$, it is more practical to use a parametrization for the functions $(\alpha_{K},\alpha_{B},\alpha_{M})$. Since we have already imposed $\alpha_{T}=0$, and our choice of $F(\phi)=1$ leads to $\alpha_{M}=0$, this parametrization allows us to separate the background evolution from the perturbation dynamics. Therefore, we adopt a conservative approach for the background assuming a standard $\Lambda$CDM evolution while the perturbations are described using the following parametrization [28]
@@
For our analysis, we take $c_{K}=1$ and $c_{B}=0.5$ and evolve the system using hi_class [32, 33, 34]. As we have mentioned the background evolution follows the standard $\Lambda$CDM model with cosmological parameters taken from [35]. Figure 1 shows the evolution of the density parameters and the linear gravitational constants associated with matter growth (34) and lensing (35).
@@ 10 Numerical analysis
In the previous section, we have constructed the basic elements required to solve (82). Starting with an initial guess for $M(k)$, the right-hand side of (82) can be integrated to yield a new function $M(k)$ until reaching convergence. We have considered four different approaches to solve this problem that we will develop in the next sections because that justifies our final approach which differs from [15]. For all our attempts, our initial guess of $M(k)$ is derived from the spherical approximation (67) which gives us
$M^{(\text{spherical})}(k)=\frac{1}{\sqrt{1+\frac{8}{3}g\Delta(k)}}$ (102)
@@
$\mathrm{B}_{\text{tree}}\left({p}_{1},{p}_{2},{p}_{3}\right)=2F_{2}\left({p}_{1},{p}_{2},{p}_{3}\right){P}(p_{1}){P}({p}_{2})+\text{ cyclic permutations, }$, (103)
@@ 10.1 Method 1
In this section, we intend to solve directly eq.(82) by assuming our two approaches for the bispectrum, A1 and A2. The integrals are performed using CUBA [38] in particular the routine Vegas but other routines have been also used as a consistency check. The results are shown in Figure (3). We see that no convergence is achieved. The function $M(k)$ is more negative at each iteration. If we increase the precision in the integral, $M(k)$ continues to go negative but less rapidly. We conclude that the system is extremely sensitive to the precision of the integral and only an infinite precision could generate the right convergence. This approach is therefore not viable and that explains why we adopt the other approaches.
Figure 3: We show the iterations on $M(k)$ at $z=0$ for the A1 bispectrum on the left and A2 bispectrum on the right. The green, red, and blue curves correspond to the initial spherical guess $M^{(0)}$, the first iteration $M^{(1)}$, and the second iteration, respectively.
@@ 10.2 Method 2
$M(k)=\frac{2\pi}{P(k)}\int_{0}^{\infty}r^{2}J_{0}(kr)\frac{\sqrt{1+4gQ_{\text{eff}}(r)\xi(r)}-1}{gQ_{\text{eff}}(r)}\,{\textnormal{d}}r,$ (111)
@@
    Starting with an initial $M(k)$, compute $\xi(r)$ and $\chi(r)$ using for example FFTLog [39].
@@
    Calculate the updated $M(k)$ using eq. (111) via another FFTLog.
After completing step 3, the new $M(k)$ is used to restart from step 1, iterating this procedure until $M(k)$ converges. The primary computational effort lies in evaluating the integral in eq.(108) which we can reduce to a 3D integral. For that we perform first a change of variables, passing from an integral over $(p_{1},p_{2})$ to an integral over $(p_{1},p_{3})$ using $\mathbf{p}_{1}+\mathbf{p}_{2}+\mathbf{p}_{3}=\mathbf{0}$
@@
The Bessel function causes oscillations in $M(k)$. However, when we split the integration into two steps, first calculating the 2D integral 117 and then applying an FFTLog over $p_{3}$ 118 gives a much better result. In contrast, directly evaluating the 3D integral 115 without using FFTLog produces large oscillations, especially at larger scales.
All methods, however, exhibit some oscillatory behavior, and in order to reduce them, we have chosen to compute these integrals for small scales, $r$, and rely on our initial guess for large scales. The outcomes for both bispectra, A1 and A2, are shown in Fig.4. After just one iteration, there is a significant deviation at large scales, where no change is expected, and the oscillations at small scales become increasingly amplified with each iteration. While it would be possible to smooth $M(k)$ for the subsequent steps, doing so would be overly artificial.
Figure 4: We show the first iteration on $M(k)$ at $z=0$ for the A1 bispectrum on the left and A2 bispectrum on the right. The green and blue curves correspond to the initial spherical guess $M^{(0)}$ and the first iteration $M^{(1)}$ respectively.
@@ 10.3 Method 3
$Q_{\text{eff}}(r)\chi^{2}(r)=\frac{1}{4\pi^{4}}\sum_{\ell=0}^{\infty}(-1)^{\ell}\iint j_{\ell}(k_{1}r)j_{\ell}(k_{2}r)M(k_{1})M(k_{2})B_{\ell}(k_{1},k_{2})k_{1}^{2}k_{2}^{2}{\textnormal{d}}k_{1}{\textnormal{d}}k_{2},$ (119)
@@
Figure 5: We show the first iteration on $M(k)$ at $z=0$ for the A1 bispectrum on the left and A2 bispectrum on the right. The green and blue curves correspond to the initial spherical guess $M^{(0)}$ and the first iteration $M^{(1)}$ respectively.
@@ 10.4 Method 4
By construction, the negative root is negative and will evolve as the solution obtained iteratively in the method 1. On the contrary, the positive root remains positive and bounded from above by $1$. This approach transforms the iteration into a quadratic relationship that enforces positivity. In conclusion, we solve for a given $M(k)$ the RHS of eq.(82) which gives us $F(k)$ once divided by $M(k)^{2}$. We therefore obtain the new $M(k)$ by the formula
@@
We found with this method a convergence with only 2 iterations. Once $M(k)$ is obtained we have the renormalized gravitational constant using eq.(77). The iteration is seen in Figure (6). We see that convergence is rapidly obtained without oscillations. Also, the form of $G_{\text{eff}}^{\text{NL}}$ is independent of the bispectrum until $k=0.1$. The slight difference at smaller scale should disappear once corrections to the bispectrum and principally to the power spectrum are taken into account.

### v2

@@ 1 Introduction
[…]
@@
Early investigations of how Vainshtein screening impacts structure formation were carried out, for instance, in the DGP scenario [15], where a non-linear perturbative approach was developed suited to scales where screening becomes important. In [16], the authors improved their N-body code using the non-linear effective gravitational constant of [15]. Our paper generalizes the methods of [15] to a class of Horndeski-like models endowed with Vainshtein screening. Specifically, we formulate a renormalized perturbation theory framework designed to capture both the large-scale, linearized modifications to gravity and the small-scale screening behavior in a self-consistent way.
@@ 2 Viable Horndeski’s theories
where $\phi$ is the scalar field and $X=-(\partial_{\mu}\phi)^{2}/2$ is the kinetic term. The functions $K$ and $G$ are arbitrary. This model corresponds to the generalized Kinetic Gravity Braiding introduced in [21].
@@ 4 Averaged effective gravitational constant
$\Phi(t,\mathbf{p})$ $=\int{\textnormal{d}}^{3}\mathbf{p}^{\prime}\left\langle\frac{\mathcal{D}\Phi(t,\boldsymbol{p})}{\mathcal{D}\delta_{m}(t,\boldsymbol{p}^{\prime})}\right\rangle\delta_{m}(t,\boldsymbol{p}^{\prime})+\frac{1}{2!}\int{\textnormal{d}}^{3}\mathbf{p}_{1}{\textnormal{d}}^{3}\mathbf{p}_{2}\left\langle\frac{\mathcal{D}^{2}\Phi(t,\boldsymbol{p})}{\mathcal{D}\delta_{m}(t,\boldsymbol{p}_{1})\mathcal{D}\delta_{m}(t,\boldsymbol{p}_{2})}\right\rangle\delta_{m}(t,\boldsymbol{p}_{1})\delta_{m}(t,\boldsymbol{p}_{2})$
$+\frac{1}{3!}\int{\textnormal{d}}^{3}\mathbf{p}_{1}{\textnormal{d}}^{3}\mathbf{p}_{2}{\textnormal{d}}^{3}\mathbf{p}_{3}\,\left\langle\frac{\mathcal{D}^{3}\Phi(t,\boldsymbol{p})}{\mathcal{D}\delta_{m}(t,\boldsymbol{p}_{1})\mathcal{D}\delta_{m}(t,\boldsymbol{p}_{2})\delta_{m}(t,\boldsymbol{p}_{3})}\right\rangle\Bigl{(}\delta_{m}(t,\mathbf{p}_{1})\delta_{m}(t,\mathbf{p}_{2})\delta_{m}(t,\mathbf{p}_{3})$
$-\langle\delta_{m}\delta_{m}\rangle(t,p_{1},p_{2})\,\delta_{m}(t,\mathbf{p}_{3})-\text{cyc.}\Bigr{)}+\ldots$ (39)
where $\mathcal{D}$ stands for a functional derivative, while $\delta_{m}(t,\mathbf{p})$ stands for the fully non-linear density contrast, and all functional derivatives are taken with respect to this same field. To prevent loop momenta already resummed inside $\delta_{m}$ from being counted a second time, we retain only the connected part of every derivative, in the sense of the cumulant (linked-cluster) expansion [15]. The square bracket in the $n=3$ term of Eq. (4) represents the connected three-point cumulant of the field $\delta_{m}$. Without the subtractions proportional to $\langle\delta_{m}\delta_{m}\rangle\,\delta_{m}$ the first product $\delta_{m}(\mathbf{p}_{1})\delta_{m}(\mathbf{p}_{2})\delta_{m}(\mathbf{p}_{3})$ would, upon Wick contraction, generate diagrams in which two external legs form a self-contained “bubble” loop; the loop momentum carried by that bubble has already been integrated over at lower (second) order. Removing these disconnected pieces therefore guarantees that every loop momentum appears exactly once in the perturbative hierarchy, preserving a one-to-one correspondence between diagrams and terms in the functional expansion. Because of translation invariance, we should have
@@ 7 Master equation
Going back to our problem, using the redefinition (76) we obtain
$\Phi(t,\mathbf{p})=\Gamma^{(1)}(t,p)\delta_{m}(t,\mathbf{p})=-\frac{1}{p^{2}}\frac{3}{2}a^{2}H^{2}\Omega_{m}\Bigl{(}1+(\mu F-1)M({t,p})\Bigr{)}\delta_{m}(t,\mathbf{p})$ (78)
from which we derive the identities
$p^{2}\Phi+\frac{3}{2}a^{2}H^{2}\Omega_{m}\delta_{m}=-\frac{3}{2}a^{2}H^{2}\Omega_{m}(\mu F-1)M\delta_{m}$ (79)
$p^{2}\Phi+\frac{3}{2}a^{2}H^{2}\Omega_{m}\mu F\delta_{m}=-\frac{3}{2}a^{2}H^{2}\Omega_{m}(\mu F-1)(M-1)\delta_{m}$ (80)
Replacing these expressions into the nonlinear Poisson equation (3), we obtain
@@ 8 Renormalized power spectrum and bispectrum
The equations (89, 90) can be solved perturbatively to correct the assumed power spectrum and bispectrum. The algorithm is the following
@@
    Step 1: Start with an initial power spectrum $P(p)$ and bispectrum $B\left(p_{1},p_{2},p_{3}\right)$.
@@
    Step 3: Insert this back into the fluid equations (89,90,91) to obtain corrected $P(p)$ and $B\left(p_{1},p_{2},p_{3}\right)$.
@@
and $D_{1}=D_{0}+D_{2}$. The equation for $D_{1}$ is therefore
$D_{1}^{\prime\prime}+\Bigl{(}2+\frac{\dot{H}}{H^{2}}\Bigr{)}D_{1}^{\prime}-\frac{3}{2}\Omega_{m}F\frac{G_{\text{eff}}^{\text{NL}}}{G}D_{1}=\frac{3}{2}\Omega_{m}F\frac{G_{\text{eff}}^{\text{NL}}}{G}D_{\text{renorm}}^{2}+2D_{\text{renorm}}^{\prime 2}$ (104)
In the squeezed limit $p_{1}\ll p_{2}$, the only piece that generates the soft pole $\propto(p_{2}/p_{1})\,\mu_{\theta}$ with $\mu_{\theta}=\hat{\mathbf{p}}_{1}\!\cdot\!\hat{\mathbf{p}}_{2}$ is the $D_{1}$ term, giving
$F_{2}^{\text{soft }}(p_{1},p_{2})=\frac{p_{2}}{2p_{1}}\mu_{\theta}D_{1}(t,k)+\text{ finite }$ (105)
[…]
$F_{2}^{\mathrm{soft}}(p_{1},p_{2})=\frac{p_{2}}{2p_{1}}\mu_{\theta}\left[D_{\text{renorm }}(t,k)\right]^{2}+\text{ finite. }$ (106)
so comparing the two expressions enforces
$D_{1}(t,k)=\left[D_{\text{renorm }}(t,k)\right]^{2}\,.$ (107)
This result can be easily recovered. Indeed, if we define the differential operator
$\mathcal{L}[X]=X^{\prime\prime}+\Bigl{(}2+\frac{\dot{H}}{H^{2}}\Bigr{)}X^{\prime}-\frac{3}{2}\Omega_{m}F\frac{G_{\text{eff}}^{\text{NL}}}{G}X$ (108)
we have $\mathcal{L}[D_{\text{renorm}}]=0$ but more interestingly
$\mathcal{L}\Bigl{[}D_{\text{renorm}}^{2}\Bigr{]}=\frac{3}{2}\Omega_{m}F\frac{G_{\text{eff}}^{\text{NL}}}{G}D_{\text{renorm}}^{2}+2D_{\text{renorm}}^{\prime 2}$ (109)
hence
$\mathcal{L}[D_{1}]=\mathcal{L}\Bigl{[}D_{\text{renorm}}^{2}\Bigr{]}$ (110)
from which we conclude
$D_{1}(t,p)=c(p)D_{\text{renorm}}(t,p)+D_{\text{renorm}}^{2}(t,p)$ (111)
where $c(p)$ is an integration constant.
For Gaussian adiabatic initial conditions (IC), at the initial epoch the connected bispectrum must vanish, so $\delta^{(2)}$ must start at order $D^{2}$. Therefore the IC are
$D_{1}\left(t_{i},p\right)=0,\quad D_{1}^{\prime}\left(t_{i},p\right)=0,$ (112)
which kills the homogeneous piece and gives
$D_{1}=D_{\text{renorm }}^{2}.$ (113)
Therefore the second-order kernel satisfies the EP/extended-Galilean soft-limit constraint; deviations from GR appear only in the finite (monopole/quadrupole) coefficients through the modified linear growth.
In conclusion, given a renormalized effective gravitational constant, obtained at previous steps, we can in principle evolve easily eqs.(95,102,103) and consequently obtain the corrections to the power spectrum and the bispectrum which will be used in eq.(85) to correct the gravitational constant. These additional steps will be performed in a future work while we will address the construction of the renormalized gravitational constant in the next sections.
@@ 9 The model
[…]
@@
[…]
@@ 10 Numerical analysis
In the previous section, we have constructed the basic elements required to solve (85). Starting with an initial guess for $M(p)$, the right-hand side of (85) can be integrated to yield a new function $M(p)$ until reaching convergence. We have considered four different approaches to solve this problem that we will develop in the next sections because that justifies our final approach which differs from [15]. For all our attempts, our initial guess of $M(p)$ is derived from the spherical approximation (67) which gives us
$M^{(\text{spherical})}(p)=\frac{1}{\sqrt{1+\frac{8}{3}g\Delta(p)}}$ (115)
@@
$\mathrm{B}_{\text{tree}}\left({p}_{1},{p}_{2},{p}_{3}\right)=2F_{2}\left({p}_{1},{p}_{2},{p}_{3}\right){P}(p_{1}){P}({p}_{2})+\text{ cyclic permutations, }$ (116)
@@ 10.1 Method 1
In this section, we intend to solve directly eq.(85) by assuming our two approaches for the bispectrum, A1 and A2. The integrals are performed using CUBA [42] in particular the routine Vegas but other routines have been also used as a consistency check. The results are shown in Figure (3). We see that no convergence is achieved. The function $M(p)$ is more negative at each iteration. If we increase the precision in the integral, $M(p)$ continues to go negative but less rapidly. We conclude that the system is extremely sensitive to the precision of the integral and only an infinite precision could generate the right convergence. This approach is therefore not viable and that explains why we adopt the other approaches.
Figure 3: We show the iterations on $M(p)$ at $z=0$ for the A1 bispectrum on the left and A2 bispectrum on the right. The green, red, and blue curves correspond to the initial spherical guess $M^{(0)}$, the first iteration $M^{(1)}$, and the second iteration, respectively.
@@ 10.2 Method 2
$M(p)=\frac{2\pi}{P(p)}\int_{0}^{\infty}r^{2}J_{0}(pr)\frac{\sqrt{1+4gQ_{\text{eff}}(r)\xi(r)}-1}{gQ_{\text{eff}}(r)}\,{\textnormal{d}}r,$ (124)
@@
    Starting with an initial $M(p)$, compute $\xi(r)$ and $\chi(r)$ using for example FFTLog [43].
@@
    Calculate the updated $M(p)$ using eq. (124) via another FFTLog.
After completing step 3, the new $M(p)$ is used to restart from step 1, iterating this procedure until $M(p)$ converges. The primary computational effort lies in evaluating the integral in eq.(121) which we can reduce to a 3D integral. For that we perform first a change of variables, passing from an integral over $(p_{1},p_{2})$ to an integral over $(p_{1},p_{3})$ using $\mathbf{p}_{1}+\mathbf{p}_{2}+\mathbf{p}_{3}=\mathbf{0}$
@@
The Bessel function causes oscillations in $M(p)$. However, when we split the integration into two steps, first calculating the 2D integral 130 and then applying an FFTLog over $p_{3}$ 131 gives a much better result. In contrast, directly evaluating the 3D integral 128 without using FFTLog produces large oscillations, especially at larger scales.
All methods, however, exhibit some oscillatory behavior, and in order to reduce them, we have chosen to compute these integrals for small scales, $r$, and rely on our initial guess for large scales. The outcomes for both bispectra, A1 and A2, are shown in Fig.4. After just one iteration, there is a significant deviation at large scales, where no change is expected, and the oscillations at small scales become increasingly amplified with each iteration. While it would be possible to smooth $M(p)$ for the subsequent steps, doing so would be overly artificial.
Figure 4: We show the first iteration on $M(p)$ at $z=0$ for the A1 bispectrum on the left and A2 bispectrum on the right. The green and blue curves correspond to the initial spherical guess $M^{(0)}$ and the first iteration $M^{(1)}$ respectively.
@@ 10.3 Method 3
$Q_{\text{eff}}(r)\chi^{2}(r)=\frac{1}{4\pi^{4}}\sum_{\ell=0}^{\infty}(-1)^{\ell}\iint j_{\ell}(p_{1}r)j_{\ell}(p_{2}r)M(p_{1})M(p_{2})B_{\ell}(p_{1},p_{2})p_{1}^{2}p_{2}^{2}{\textnormal{d}}p_{1}{\textnormal{d}}p_{2},$ (132)
@@
Figure 5: We show the first iteration on $M(p)$ at $z=0$ for the A1 bispectrum on the left and A2 bispectrum on the right. The green and blue curves correspond to the initial spherical guess $M^{(0)}$ and the first iteration $M^{(1)}$ respectively.
@@ 10.4 Method 4
By construction, the negative root is negative and will evolve as the solution obtained iteratively in the method 1. On the contrary, the positive root remains positive and bounded from above by $1$. This approach transforms the iteration into a quadratic relationship that enforces positivity. In conclusion, we solve for a given $M(p)$ the RHS of eq.(85) which gives us $F(k)$ once divided by $M(p)^{2}$. We therefore obtain the new $M(p)$ by the formula
@@
We found with this method a convergence with only 2 iterations. Once $M(p)$ is obtained we have the renormalized gravitational constant using eq.(77). The iteration is seen in Figure (6). We see that convergence is rapidly obtained without oscillations. Also, the form of $G_{\text{eff}}^{\text{NL}}$ is independent of the bispectrum until $k=0.1$. The slight difference at smaller scale should disappear once corrections to the bispectrum and principally to the power spectrum are taken into account.
@@
11 Comparison with Standard Perturbation Theory at 1-Loop
In the previous sections, we developed the formalism of the renormalized approach to the nonlinear effective gravitational constant, which reduces to GR at small scales. This formalism allows us to identify the scale at which the Vainshtein mechanism becomes effective. To assess the advantages of this approach, we now compare it with the standard perturbation theory (SPT) by performing the same calculations within the SPT framework.
In SPT, all perturbation fields, $\Phi$, $\theta$, $\delta_{m}$, and $Q$, are expanded as
$Y(t,\mathbf{p})=\sum_{n=1}Y_{n}(t,\mathbf{p})$ (141)
where $Y$ stands for any of the perturbation fields, and $Y_{n}$ represents the $n$th-order term in the expansion. At linear order, one finds
$\delta_{m}^{(1)}(t,\mathbf{p})=D_{+}(t)\delta_{1}(\mathbf{p})\,,\qquad-p^{2}\Phi_{1}(t,\mathbf{p})=\frac{3}{2}a^{2}H^{2}\Omega_{m}\mu FD_{+}(t)\delta_{\mathrm{L}}(\mathbf{p}),$ (142)
where $D_{+}(t)$ is the growing mode of the linear evolution equation (95)
$D^{\prime\prime}+\Bigl{(}2+\frac{\dot{H}}{H^{2}}\Bigr{)}D^{\prime}-\frac{3}{2}\Omega_{m}F\mu D=0$ (143)
with primes denoting derivatives with respect to $\ln a$. The second-order solution $(\delta_{m}^{(2)}(t,\mathbf{p}),\Phi_{2}(t,\mathbf{p}))$ has been derived in [45], and the third-order $(\delta_{m}^{(3)}(t,\mathbf{p}),\Phi_{3}(t,\mathbf{p}))$ in [46]. Knowing this expansion we can obtain the effective gravitational constant. Indeed, from Eq.(71), we have
$-p^{2}\Phi(t,\mathbf{p})=\frac{3}{2}a^{2}H^{2}\Omega_{m}F\frac{G_{\text{eff}}^{\text{NL}}}{G}\delta_{m}(t,\mathbf{p})$ (144)
contracting both sides with $\delta_{m}(t,-\mathbf{p})$ and taking an ensemble average, we obtain
$-p^{2}P_{\Phi\delta}=\frac{3}{2}a^{2}H^{2}\Omega_{m}F\frac{G_{\text{eff}}^{\text{NL}}}{G}P_{\delta\delta}$ (145)
from which we get
$\frac{G_{\text{eff}}^{\text{NL}}(t,{p})}{G}=-\frac{p^{2}}{\frac{3}{2}a^{2}H^{2}\Omega_{m}F}\frac{P_{\Phi\delta}(t,{p})}{P_{\delta\delta}(t,{p})}$ (146)
where we have defined the cross correlation function
$\Bigl{\langle}\Phi(\mathbf{p})\delta_{m}(\mathbf{k})\Bigr{\rangle}$ $=(2\pi)^{3}\delta_{D}(\mathbf{p}+\mathbf{k})P_{\Phi\delta}(p)$ (147)
and $P_{\delta\delta}\equiv P(k)$. Details of the calculations are provided in Appendix B. At 1-loop order, we obtain from Eq. (146) the following expression
$\frac{G_{\text{eff}}^{\text{NL}}}{G}$ $\simeq-\frac{p^{2}}{\frac{3}{2}a^{2}H^{2}\Omega_{m}F}\left(\frac{P_{\Phi\delta}^{\text{lin.}}}{P_{L}}+\frac{P_{\Phi\delta}^{\text{1-loop}}}{P_{L}}-\frac{P_{\Phi\delta}^{\text{lin.}}}{P_{L}}\frac{P_{\delta\delta}^{\text{1-loop}}}{P_{L}}\right)$ (148)
$=\mu-\frac{p^{2}}{\frac{3}{2}a^{2}H^{2}\Omega_{m}F}\frac{P_{\Phi\delta}^{\text{1-loop}}}{P_{L}}-\mu\frac{P_{\delta\delta}^{\text{1-loop}}}{P_{L}}$ (149)
where we defined
$\langle\delta_{1}(t,\mathbf{p})\delta_{1}(t,\mathbf{k})\rangle=(2\pi)^{3}\delta_{D}(\mathbf{p}+\mathbf{k})P_{L}(p)$ (150)
The numerical results are shown in Fig. (8), which illustrates that the 1-loop correction does not fully capture the behavior of the gravitational constant in the non-linear regime.
Figure 8: We show the gravitational constant at $z=0$. The blue line indicates the renormalized gravitational constant, while the red line corresponds to the SPT approach.
At very large scales and early times (the IR limit), the 1-loop power is controlled by the hard region, where the loop momentum $q$ is finite so that $p\rightarrow 0$ implies $r=q/p\rightarrow\infty$. The soft region $q\ll p$ is subleading. One finds
$P_{\delta\delta}^{\text{1-loop}}(p\rightarrow 0)$ $=2P_{\delta\delta}^{13}(p\rightarrow 0)+\mathcal{O}(p^{4})$ (151)
$=\frac{147-144\lambda_{1}-64\lambda_{2}}{630\pi^{2}}p^{2}P_{L}(p)\int_{0}^{\infty}dqP_{L}(q)+\mathcal{O}(p^{4})$ (152)
where $\lambda_{1}$ and $\lambda_{2}$ are defined in Eqs. (204,205). Likewise,
$P_{\Phi\delta}^{\text{1-loop}}(p\rightarrow 0)$ $=P_{\Phi\delta}^{13}(p\rightarrow 0)+P_{\Phi\delta}^{31}(p\rightarrow 0)+\mathcal{O}(p^{2})$ (153)
$=-\frac{a^{2}H^{2}}{\pi^{2}}P_{L}(p)\left[\kappa_{\Phi}\left(\frac{7}{30}-\frac{8}{35}\lambda_{1}-\frac{32}{315}\lambda_{2}\right)+\frac{8}{15}\left(\sigma_{\Phi}+\mu_{\Phi}\right)\right]\int_{0}^{\infty}dqP_{L}(q)+\mathcal{O}(p^{2})$ (154)
with $(\sigma_{\Phi},\mu_{\Phi})$ given in Eqs. (189,190). Hence, in the IR,
$\frac{G_{\text{eff}}^{\text{NL}}}{G}$ $=\mu+\left[\frac{1}{\Omega_{m}F}\left(\kappa_{\Phi}\left(\frac{7}{45}-\frac{16}{105}\lambda_{1}-\frac{64}{945}\lambda_{2}\right)+\frac{16}{45}\left(\sigma_{\Phi}+\mu_{\Phi}\right)\right)\right$.
$\qquad\qquad\qquad\qquad\left.-\mu\frac{147-144\lambda_{1}-64\lambda_{2}}{630}\right]\frac{p^{2}}{\pi^{2}}\int_{0}^{\infty}dqP_{L}(q)+\mathcal{O}(p^{4})$ (155)
We can check that this matches the expected GR matter-era limit: taking $\mu=\Omega_{m}=F=\lambda_{1}=\lambda_{2}=1$, $\kappa_{\Phi}=3/2$ and $\sigma_{\Phi}=\mu_{\Phi}=0$, all coefficients cancel and
$\frac{G_{\text{eff}}^{\text{NL}}}{G}=1+\mathcal{O}(p^{4})$ (156)
In summary, at very large scales in the early Universe the one-loop correction scales as $p^{2}$. Assuming matter domination, the coefficients in (11) follow from the time dependences of the growing and decaying modes, which during the matter era read [47, 48]
$D_{\pm}(t)=t^{p_{\pm}}\,,\quad p_{\pm}=\frac{-1\pm\sqrt{1+24F\mu}}{6}$ (157)
From these relations one obtains $\lambda_{1},\lambda_{2},\kappa_{\Phi},\sigma_{\Phi}$ and $\mu_{\Phi}$, so that (11) becomes
$\frac{G_{\text{eff}}^{\text{NL}}}{G}=\mu-Ag\frac{\mu F-1}{F}p^{2}+\mathcal{O}(p^{4})$ (158)
with
$A=\frac{32}{15\pi^{2}}\frac{X(X-1)+2g(6+X-X^{2})}{(X-1)(3X-1)}\int_{0}^{\infty}dqP_{L}(q)$ (159)
where $X=\sqrt{1+24\mu F}$.
We now verify that the renormalized approach yields the same IR behavior, as it should in the linear regime. In the IR limit (the squeezed bispectrum), the leading term is
$B(\mathbf{p}_{1},\mathbf{p}-\mathbf{p}_{1})=P_{L}(p)\left(\Bigl{(}\frac{13}{7}+\frac{8}{7}\mu_{\theta}^{2}\Bigr{)}P_{L}(p_{1})-p_{1}\mu_{\theta}^{2}P_{L}^{\prime}(p_{1})\right)+\mathcal{O}(p)$ (160)
Using (85), the leading-order behavior of $M(p)$ is
$M(p)=1-\frac{g}{(2\pi)^{2}}p^{2}\int dqP_{L}(q)M(q)^{2}\left(\frac{292}{105}-\frac{4}{15}\frac{d\ln P_{L}}{d\ln q}\right)+\mathcal{O}(p^{3})$ (161)
This equation is solved recursively until convergence. Once converged, the integral is a number, so we write
$\frac{1}{(2\pi)^{2}}\int dqP_{L}(q)M(q)^{2}\left(\frac{292}{105}-\frac{4}{15}\frac{d\ln P_{L}}{d\ln q}\right)=B$ (162)
which implies
$M(p)=1-Bgp^{2}+\mathcal{O}(p^{3})$ (163)
Remark that in the IR regime the source in the master equation is non-local: $\sim p^{2}\int dqP_{L}(q)$ (after factoring the external $P_{L}(p)$), not $\propto P_{L}(k)^{2}$. Hence a naive small-$P$ expansion $M\simeq 1-cP+\ldots$ does not apply. From (77), the renormalized effective gravitational constant in the IR limit is
$\frac{G_{\text{eff}}^{\text{NL}}}{G}=\mu-Bg\frac{\mu F-1}{F}p^{2}+\mathcal{O}(p^{3})$ (164)
The IR limits coincide if $B=A$ namely
$x\int dqP_{L}(q)M(q)^{2}\left(\frac{292}{105}-\frac{4}{15}\frac{d\ln P_{L}}{d\ln q}\right)=\frac{128}{15}\frac{X(X-1)+2g(6+X-X^{2})}{(X-1)(3X-1)}\int_{0}^{\infty}dqP_{L}(q)$ (165)
Even if the IR limit cannot be checked analytically, we conclude that the $p^{2}$ scaling is recovered in both approaches. At the next order a difference may already appear: the renormalized approach scales as $p^{3}$ (if absent cancellations), whereas the one-loop correction scales as $p^{4}$.